OpenStudy (anonymous):

Find the limit as t tends to ZERO from the RIGHT. lim 21sqrt(t) / t t->0+

OpenStudy (anonymous):

21 is NOT the answer

myininaya (myininaya):

\[\lim_{t \rightarrow 0^+}\frac{21 \sqrt{t}}{t} \cdot \frac{\sqrt{t}}{\sqrt{t}}=\lim_{t \rightarrow 0^+}\frac{21(t)}{t \sqrt{t}}=\lim_{t \rightarrow 0^+}\frac{21}{\sqrt{t}}\] Does not exist To be more exact if approaches positive infinity

myininaya (myininaya):

it*

OpenStudy (anonymous):

so any time i see a square root with a zero inside, the answer will be infinity or DNE?

OpenStudy (anonymous):

\[\sqrt{t}\] is in the denominator, so as numbers get closer and closer to 0, the limit approaches infinity